# Splittings of One-Ended Groups with One-Ended Halfspaces

### v3

@@ 4.4 The pocset $\mathcal{P}$
#=$\mathcal{H}^{\#}=\{\mathfrak{h}\in\mathcal{H}(X)\mid\hat{\mathfrak{h}}\notin G\cdot\hat{\mathfrak{h}}_{0}\},$
and let $\mathcal{M}_{0}$ denote the set of $\sim$-classes in $V\mathfrak{h}_{0}$ that are not contained in a bounded neighborhood of $\hat{\mathfrak{h}}_{0}$. Note that the $\sim$-class $[x]$ appearing in Lemma 4.13 is in $\mathcal{M}_{0}$, because the halfspace $\mathfrak{h}\subseteq[x]$ is not contained in a bounded neighborhood of $\hat{\mathfrak{h}}_{0}$. The same goes for the $\sim$-classes $[x]$ and $[y]$ appearing in Lemma 4.14. We also get the following lemma.
@@
We cannot have $[x]\preceq[x]^{*}$ for $[x]\in\mathcal{M}_{0}$.
@@
Any path from $[x]$ to $[x]^{*}$ must either pass through $\hat{\mathfrak{h}}_{0}$ or through a $G_{0}$-translate of $\delta$, which is contained in a bounded neighborhood of $\hat{\mathfrak{h}}_{0}$. The lemma follows since $[x]$ is not contained in a bounded neighborhood of $\hat{\mathfrak{h}}_{0}$ (which is by definition of $\mathcal{M}_{0}$). ∎
@@
$(\mathfrak{a},\mathfrak{h})=\begin{cases}(\mathfrak{h},\mathfrak{h}),&\mathfrak{h}\in\mathcal{H}^{\#},\\ (g[x],g\mathfrak{h}_{0}),&g\in G,\,[x]\in\mathcal{M}_{0},\\ (g[x]^{*},g\mathfrak{h}_{0}^{*}),&g\in G,\,[x]\in\mathcal{M}_{0},\end{cases}$ (8)
@@
For distinct $[x],[y]\in\mathcal{M}_{0}$, we have $([x],\mathfrak{h}_{0})\leq([y]^{*},\mathfrak{h}_{0}^{*})$, but we do not have $([x],\mathfrak{h}_{0})\leq([y],\mathfrak{h}_{0})$ or $([x]^{*},\mathfrak{h}_{0}^{*})\leq([y]^{*},\mathfrak{h}_{0}^{*})$.
@@
We know that $[x]$ and $[y]$ are disjoint, so $[x]\preceq[y]^{*}$. Moreover, $\mathfrak{h}_{0}^{*}-\hat{\mathfrak{h}}_{0}\subseteq[y]^{*}$ is not contained in a bounded neighborhood of $[x]$, so $[x]\prec[y]^{*}$ and $([x],\mathfrak{h}_{0})\leq([y]^{*},\mathfrak{h}_{0}^{*}h)^{*}$ is an inequality of type (1).
Secondly, we cannot have $([x],\mathfrak{h}_{0})\leq([y],\mathfrak{h}_{0})$ as that would imply $[x]\preceq[y]\subseteq[x]^{*}$, contradicting Lemma 4.15. Similarly, we cannot have $([x]^{*},\mathfrak{h}_{0}^{*})\leq([y]^{*},\mathfrak{h}_{0}^{*})$ as that would imply $[y]\subseteq[x]^{*}\preceq[y]^{*}$, again contradicting Lemma 4.15. ∎
@@
Reflexivity holds by (2). Let’s now prove antisymmetry. If $(\mathfrak{a}_{1},\mathfrak{h}_{1})\leq(\mathfrak{a}_{2},\mathfrak{h}_{2})\leq(\mathfrak{a}_{1},\mathfrak{h}_{1})$ then $\mathfrak{a}_{1}\preceq\mathfrak{a}_{2}\preceq\mathfrak{a}_{1}$, so both $\leq$ inequalities must be of type (2), which means that $\mathfrak{h}_{1}\subseteq\mathfrak{h}_{2}\subseteq\mathfrak{h}_{1}$, and so $\mathfrak{h}_{1}=\mathfrak{h}_{2}$. If $\mathfrak{h}_{1}\in\mathcal{H}^{\#}$ then $(\mathfrak{a}_{1},\mathfrak{h}_{1})=(\mathfrak{a}_{2},\mathfrak{h}_{2})=(\mathfrak{h}_{1},\mathfrak{h}_{1})$. If $\mathfrak{h}_{1}=g\mathfrak{h}_{0}$ (some $g\in G$), then $(\mathfrak{a}_{1},\mathfrak{h}_{1})=(g[x],g\mathfrak{h}_{0})$ and $(\mathfrak{a}_{2},\mathfrak{h}_{2})=(g[y],g\mathfrak{h}_{0})$ for some $[x],[y]\in\mathcal{M}_{0}$, and $[x]=[y]$ by Lemma 4.17. Similarly, if $\mathfrak{h}_{1}=g\mathfrak{h}_{0}^{*}$ (some $g\in G$), then $(\mathfrak{a}_{1},\mathfrak{h}_{1})=(g[x]^{*},g\mathfrak{h}_{0}^{*})$ and $(\mathfrak{a}_{2},\mathfrak{h}_{2})=(g[y]^{*},g\mathfrak{h}_{0}^{*})$ for some $[x],[y]\in\mathcal{M}_{0}$, and $[x]=[y]$ by Lemma 4.17.
@@
$(\mathfrak{h},\mathfrak{h})$ $\mapsto(\mathfrak{h}^{*},\mathfrak{h}^{*}),$ $\mathfrak{h}\in\mathcal{H}^{\#},$
@@
If both inequalities did hold, then we would have $\mathfrak{a}_{1}\preceq\mathfrak{a}_{2}$ and $\mathfrak{a}_{1}^{*}\preceq\mathfrak{a}_{2}$, so $X\preceq\mathfrak{a}_{2}$ and $\mathfrak{a}_{2}^{*}\preceq\mathfrak{a}_{2}$. If $(\mathfrak{a}_{2},\mathfrak{h}_{2})$ takes the form $(\mathfrak{h},\mathfrak{h})$ or $(g[x],g\mathfrak{h}_{0})$, then $\mathfrak{a}_{2}\subseteq\mathfrak{h}_{2}$ and we get a contradiction since $\mathfrak{h}_{2}^{*}\npreceq\mathfrak{h}_{2}$ (Lemma 3.2). If $(\mathfrak{a}_{2},\mathfrak{h}_{2})$ takes the form $(g[x]^{*},g\mathfrak{h}_{0}^{*})$, then $\mathfrak{a}_{2}^{*}=g[x]\preceq g[x]^{*}$, contradicting Lemma 4.15. ∎
@@
We will spend the rest of the proof showing that one of the rows of (9) holds. Among elements of the form $(\mathfrak{h},\mathfrak{h})$ in $\mathcal{P}$ (with $\mathfrak{h}\in\mathcal{H}^{\#}$), the relation $\leq$ is determined by the projection of these elements to the pocset $(\mathcal{H}(X),\subseteq)$. Since this pocset has no pairs of transverse elements, we deduce that if $(\mathfrak{a}_{1},\mathfrak{h}_{1}),(\mathfrak{a}_{2},\mathfrak{h}_{2})$ are both of the form $(\mathfrak{h},\mathfrak{h})$ with $\mathfrak{h}\in\mathcal{H}^{\#}$, then one of the rows of (9) holds.
Henceforth, assume that $(\mathfrak{a}_{1},\mathfrak{h}_{1}),(\mathfrak{a}_{2},\mathfrak{h}_{2})$ are not both of the form $(\mathfrak{h},\mathfrak{h})$ with $\mathfrak{h}\in\mathcal{H}^{\#}$. Without loss of generality, assume $(\mathfrak{a}_{1},\mathfrak{h}_{1})$ is not of this form, so $(\mathfrak{a}_{1},\mathfrak{h}_{1})$ is of the form $(g[x],g\mathfrak{h}_{0})$ or $(g[x]^{*},g\mathfrak{h}_{0}^{*})$. Since $G$ acts on $\mathcal{P}$ respecting $\leq$ and *, it suffices to consider the case where $(\mathfrak{a}_{1},\mathfrak{h}_{1})$ is of the form $([x],\mathfrak{h}_{0})$ or $([x]^{*},\mathfrak{h}_{0}^{*})$. And replacing $(\mathfrak{a}_{1},\mathfrak{h}_{1})$ with $(\mathfrak{a}_{1},\mathfrak{h}_{1})^{*}$ if necessary, we may assume $(\mathfrak{a}_{1},\mathfrak{h}_{1})=([x],\mathfrak{h}_{0})$. We now have three cases to consider:
@@
    Suppose $\mathfrak{h}_{2}\in\{\mathfrak{h}_{0},\mathfrak{h}_{0}^{*}\}$. Replacing $(\mathfrak{a}_{2},\mathfrak{h}_{2})$ with $(\mathfrak{a}_{2},\mathfrak{h}_{2})^{*}$ if necessary, we may assume that $\mathfrak{h}_{2}=\mathfrak{h}_{0}$. Then $(\mathfrak{a}_{2},\mathfrak{h}_{2})=([y],\mathfrak{h}_{0})$ for some $[y]\in\mathcal{M}_{0}$. Since $(\mathfrak{a}_{1},\mathfrak{h}_{1}),(\mathfrak{a}_{2},\mathfrak{h}_{2})$ are distinct, we have $[x]\neq[y]$, so the second row of (9) holds by Lemma 4.17.
@@
    Suppose $\mathfrak{h}_{2}\subsetneq\mathfrak{h}_{0}$ or $\mathfrak{h}_{2}^{*}\subsetneq\mathfrak{h}_{0}$. Replacing $(\mathfrak{a}_{2},\mathfrak{h}_{2})$ with $(\mathfrak{a}_{2},\mathfrak{h}_{2})^{*}$ if necessary, we may assume that $\mathfrak{h}_{2}\subsetneq\mathfrak{h}_{0}$. By Lemma 4.13, there is a unique $[y]\in\mathcal{M}_{0}$ with $\mathfrak{h}_{2}\subseteq[y]$. Let’s first show that either the third or fourth row of (9) holds in the case $[x]=[y]$. We split into three subcases:
@@
      Suppose $(\mathfrak{a}_{2},\mathfrak{h}_{2})=(\mathfrak{h}_{2},\mathfrak{h}_{2})$. Then $\mathfrak{a}_{2}=\mathfrak{h}_{2}\subseteq[y]=\mathfrak{a}_{1}$, so the fourth row of (9) holds (and they are inequalities of type (2)).
@@
      Suppose $(\mathfrak{a}_{2},\mathfrak{h}_{2})=(g[z],g\mathfrak{h}_{0})$, with $g\in G$, $[z]\in\mathcal{M}_{0}$. Then $\mathfrak{a}_{2}=g[z]\subseteq\mathfrak{h}_{2}\subseteq[y]=\mathfrak{a}_{1}$, so again the fourth row of (9) holds.
@@
      Suppose $(\mathfrak{a}_{2},\mathfrak{h}_{2})=(g[z]^{*},g\mathfrak{h}_{0}^{*})$, with $g\in G$, $[z]\in\mathcal{M}_{0}$. Let $[y^{\prime}]$ be the $\sim$-class with $\mathfrak{h}_{0}^{*}\subseteq g[y^{\prime}]$. By Lemma 4.14 (applied both ways round), we have $g[y^{\prime}]^{*}\preceq[y]$ and $[y]^{*}\preceq g[y^{\prime}]$, so
@@
      If $[z]=[y^{\prime}]$, then the fourth row of (9) holds by (10). If $[z]\neq[y^{\prime}]$, then
@@
      If $(\mathfrak{a}_{2},\mathfrak{h}_{2})=(\mathfrak{h}_{2},\mathfrak{h}_{2})$, then $[x]\subseteq\mathfrak{h}_{0}\subseteq\mathfrak{h}_{2}$, so the first row of (9) holds (and they are both inequalities of type (2)).
@@
      If $(\mathfrak{a}_{2},\mathfrak{h}_{2})=(g[z],g\mathfrak{h}_{0})$, with $g\in G$ and $[z]\in\mathcal{M}_{0}$, then up to the action of $G$ (and with the roles of $(\mathfrak{a}_{1},\mathfrak{h}_{1}),(\mathfrak{a}_{2},\mathfrak{h}_{2})$ reversed) this is the same as case (2).

### v4

@@ 4.6 The pocset $\mathcal{P}$
$\mathcal{H}^{\#}=\{\mathfrak{h}\in\mathcal{H}(X)\mid\hat{\mathfrak{h}}\notin G\cdot\hat{\mathfrak{h}}_{0}\},$
and let $\mathcal{M}_{0}$ denote the set of $\sim$-classes in $V\mathfrak{h}_{0}$ that are not contained in a bounded neighborhood of $\hat{\mathfrak{h}}_{0}$. Note that the $\sim$-class $[x]$ appearing in Lemma 4.14 is in $\mathcal{M}_{0}$, because the halfspace $V\mathfrak{h}\subseteq[x]$ is not contained in a bounded neighborhood of $\hat{\mathfrak{h}}_{0}$. The same goes for the $\sim$-classes $[x]$ and $[y]$ appearing in Lemma 4.15. We also get the following lemma.
@@
We cannot have $x]\preceq[x]^{*}$ for $x]\in\mathcal{M}_{0}$.
@@
Any path from $[x]$ to $x]^{*}$ must either pass through $\hat{\mathfrak{h}}_{0}$ or through a $G_{0}$-translate of $\delta$, which is contained in a bounded neighborhood of $\hat{\mathfrak{h}}_{0}$. The lemma follows since $[x]$ is not contained in a bounded neighborhood of $\hat{\mathfrak{h}}_{0}$ (which is by definition of $\mathcal{M}_{0}$). ∎
@@
$(\mathfrak{a},\mathfrak{h})=\begin{cases}(V\mathfrak{h},\mathfrak{h}),&\mathfrak{h}\in\mathcal{H}^{\#},\\ (g[x],g\mathfrak{h}_{0}),&g\in G,\,[x]\in\mathcal{M}_{0},\\ (g[x]^{*},g\mathfrak{h}_{0}^{*}),&g\in G,\,[x]\in\mathcal{M}_{0},\end{cases}$ (11)
@@
For distinct $x],[y]\in\mathcal{M}_{0}$, we have $([x],\mathfrak{h}_{0})\leq([y]^{*},\mathfrak{h}_{0}^{*})$, but we do not have $([x],\mathfrak{h}_{0})\leq([y],\mathfrak{h}_{0})$ or $([x]^{*},\mathfrak{h}_{0}^{*})\leq([y]^{*},\mathfrak{h}_{0}^{*})$.
@@
We know that $[x]$ and $[y]$ are disjoint, so $x]\preceq[y]^{*}$. Moreover, $V\mathfrak{h}_{0}^{*}-\hat{\mathfrak{h}}_{0}\subseteq[y]^{*}$ is not contained in a bounded neighborhood of $[x]$, so $x]\prec[y]^{*}$ and $([x],\mathfrak{h}_{0})\leq([y]^{*},\mathfrak{h}_{0}^{*}h)^{*}$ is an inequality of type (1).
Secondly, we cannot have $([x],\mathfrak{h}_{0})\leq([y],\mathfrak{h}_{0})$ as that would imply $x]\preceq[y]\subseteq[x]^{*}$, contradicting Lemma 4.16. Similarly, we cannot have $([x]^{*},\mathfrak{h}_{0}^{*})\leq([y]^{*},\mathfrak{h}_{0}^{*})$ as that would imply $y]\subseteq[x]^{*}\preceq[y]^{*}$, again contradicting Lemma 4.16. ∎
@@
Reflexivity holds by (2). Let’s now prove antisymmetry. If $(\mathfrak{a}_{1},\mathfrak{h}_{1})\leq(\mathfrak{a}_{2},\mathfrak{h}_{2})\leq(\mathfrak{a}_{1},\mathfrak{h}_{1})$ then $\mathfrak{a}_{1}\preceq\mathfrak{a}_{2}\preceq\mathfrak{a}_{1}$, so both $\leq$ inequalities must be of type (2), which means that $\mathfrak{h}_{1}\subseteq\mathfrak{h}_{2}\subseteq\mathfrak{h}_{1}$, and so $\mathfrak{h}_{1}=\mathfrak{h}_{2}$. If $\mathfrak{h}_{1}\in\mathcal{H}^{\#}$ then $(\mathfrak{a}_{1},\mathfrak{h}_{1})=(\mathfrak{a}_{2},\mathfrak{h}_{2})=(V\mathfrak{h}_{1},\mathfrak{h}_{1})$. If $\mathfrak{h}_{1}=g\mathfrak{h}_{0}$ (some $g\in G$), then $(\mathfrak{a}_{1},\mathfrak{h}_{1})=(g[x],g\mathfrak{h}_{0})$ and $(\mathfrak{a}_{2},\mathfrak{h}_{2})=(g[y],g\mathfrak{h}_{0})$ for some $x],[y]\in\mathcal{M}_{0}$, and $[x]=[y]$ by Lemma 4.18. Similarly, if $\mathfrak{h}_{1}=g\mathfrak{h}_{0}^{*}$ (some $g\in G$), then $(\mathfrak{a}_{1},\mathfrak{h}_{1})=(g[x]^{*},g\mathfrak{h}_{0}^{*})$ and $(\mathfrak{a}_{2},\mathfrak{h}_{2})=(g[y]^{*},g\mathfrak{h}_{0}^{*})$ for some $x],[y]\in\mathcal{M}_{0}$, and $[x]=[y]$ by Lemma 4.18.
@@
$(V\mathfrak{h},\mathfrak{h})$ $\mapsto(V\mathfrak{h}^{*},\mathfrak{h}^{*}),$ $\mathfrak{h}\in\mathcal{H}^{\#},$
@@
Write $(\mathfrak{a}_{1},\mathfrak{h}_{1})^{*}=(\mathfrak{a}_{1}^{*},\mathfrak{h}_{1}^{*})$. If both inequalities in the lemma did hold, then we would have $\mathfrak{a}_{1}\preceq\mathfrak{a}_{2}$ and $\mathfrak{a}_{1}^{*}\preceq\mathfrak{a}_{2}$, so $X\preceq\mathfrak{a}_{2}$ and $\mathfrak{a}_{2}^{*}\preceq\mathfrak{a}_{2}$. If $(\mathfrak{a}_{2},\mathfrak{h}_{2})$ takes the form $(V\mathfrak{h},\mathfrak{h})$ or $(g[x],g\mathfrak{h}_{0})$, then $\mathfrak{a}_{2}\subseteq\mathfrak{h}_{2}$ and we get a contradiction since $\mathfrak{h}_{2}^{*}\npreceq\mathfrak{h}_{2}$ (Lemma 3.3). If $(\mathfrak{a}_{2},\mathfrak{h}_{2})$ takes the form $(g[x]^{*},g\mathfrak{h}_{0}^{*})$, then $\mathfrak{a}_{2}^{*}=g[x]\preceq g[x]^{*}$, contradicting Lemma 4.16. ∎
@@
We will spend the rest of the proof showing that one of the rows of (12) holds. Among elements of the form $(V\mathfrak{h},\mathfrak{h})$ in $\mathcal{P}$ (with $\mathfrak{h}\in\mathcal{H}^{\#}$), the relation $\leq$ is determined by the projection of these elements to the pocset $(\mathcal{H}(X),\subseteq)$. Since this pocset has no pairs of transverse elements, we deduce that if $(\mathfrak{a}_{1},\mathfrak{h}_{1}),(\mathfrak{a}_{2},\mathfrak{h}_{2})$ are both of the form $(V\mathfrak{h},\mathfrak{h})$ with $\mathfrak{h}\in\mathcal{H}^{\#}$, then one of the rows of (12) holds.
Henceforth, assume that $(\mathfrak{a}_{1},\mathfrak{h}_{1}),(\mathfrak{a}_{2},\mathfrak{h}_{2})$ are not both of the form $(V\mathfrak{h},\mathfrak{h})$ with $\mathfrak{h}\in\mathcal{H}^{\#}$. Without loss of generality, assume $(\mathfrak{a}_{1},\mathfrak{h}_{1})$ is not of this form, so $(\mathfrak{a}_{1},\mathfrak{h}_{1})$ is of the form $(g[x],g\mathfrak{h}_{0})$ or $(g[x]^{*},g\mathfrak{h}_{0}^{*})$. Since $G$ acts on $\mathcal{P}$ respecting $\leq$ and *, it suffices to consider the case where $(\mathfrak{a}_{1},\mathfrak{h}_{1})$ is of the form $([x],\mathfrak{h}_{0})$ or $([x]^{*},\mathfrak{h}_{0}^{*})$. And replacing $(\mathfrak{a}_{1},\mathfrak{h}_{1})$ with $(\mathfrak{a}_{1},\mathfrak{h}_{1})^{*}$ if necessary, we may assume $(\mathfrak{a}_{1},\mathfrak{h}_{1})=([x],\mathfrak{h}_{0})$. We now have three cases to consider:
@@
    Suppose $\mathfrak{h}_{2}\in\{\mathfrak{h}_{0},\mathfrak{h}_{0}^{*}\}$. Replacing $(\mathfrak{a}_{2},\mathfrak{h}_{2})$ with $(\mathfrak{a}_{2},\mathfrak{h}_{2})^{*}$ if necessary, we may assume that $\mathfrak{h}_{2}=\mathfrak{h}_{0}$. Then $(\mathfrak{a}_{2},\mathfrak{h}_{2})=([y],\mathfrak{h}_{0})$ for some $y]\in\mathcal{M}_{0}$. Since $(\mathfrak{a}_{1},\mathfrak{h}_{1}),(\mathfrak{a}_{2},\mathfrak{h}_{2})$ are distinct, we have $x]\neq[y]$, so the second row of (12) holds by Lemma 4.18.
@@
    Suppose $\mathfrak{h}_{2}\subsetneq\mathfrak{h}_{0}$ or $\mathfrak{h}_{2}^{*}\subsetneq\mathfrak{h}_{0}$. Replacing $(\mathfrak{a}_{2},\mathfrak{h}_{2})$ with $(\mathfrak{a}_{2},\mathfrak{h}_{2})^{*}$ if necessary, we may assume that $\mathfrak{h}_{2}\subsetneq\mathfrak{h}_{0}$. By Lemma 4.14, there is a unique $y]\in\mathcal{M}_{0}$ with $V\mathfrak{h}_{2}\subseteq[y]$. Let’s first show that either the third or fourth row of (12) holds in the case $[x]=[y]$. We split into three subcases:
@@
      Suppose $(\mathfrak{a}_{2},\mathfrak{h}_{2})=(V\mathfrak{h}_{2},\mathfrak{h}_{2})$. Then $\mathfrak{a}_{2}=V\mathfrak{h}_{2}\subseteq[y]=\mathfrak{a}_{1}$, so the fourth row of (12) holds (and they are inequalities of type (2)).
@@
      Suppose $(\mathfrak{a}_{2},\mathfrak{h}_{2})=(g[z],g\mathfrak{h}_{0})$, with $g\in G$, $z]\in\mathcal{M}_{0}$. Then $\mathfrak{a}_{2}=g[z]\subseteq V\mathfrak{h}_{2}\subseteq[y]=\mathfrak{a}_{1}$, so again the fourth row of (12) holds.
@@
      Suppose $(\mathfrak{a}_{2},\mathfrak{h}_{2})=(g[z]^{*},g\mathfrak{h}_{0}^{*})$, with $g\in G$, $z]\in\mathcal{M}_{0}$. Let $y^{\prime}]$ be the $\sim$-class with $V\mathfrak{h}_{0}^{*}\subseteq g[y^{\prime}]$. By Lemma 4.15 (applied both ways round), we have $g[y^{\prime}]^{*}\preceq[y]$ and $y]^{*}\preceq g[y^{\prime}]$, so
@@
      If $z]=[y^{\prime}]$, then the fourth row of (12) holds by (13). If $z]\neq[y^{\prime}]$, then
@@
      If $(\mathfrak{a}_{2},\mathfrak{h}_{2})=(V\mathfrak{h}_{2},\mathfrak{h}_{2})$, then $x]\subseteq\mathfrak{h}_{0}\subseteq\mathfrak{h}_{2}$, so the first row of (12) holds (and they are both inequalities of type (2)).
@@
      If $(\mathfrak{a}_{2},\mathfrak{h}_{2})=(g[z],g\mathfrak{h}_{0})$, with $g\in G$ and $z]\in\mathcal{M}_{0}$, then up to the action of $G$ (and with the roles of $(\mathfrak{a}_{1},\mathfrak{h}_{1}),(\mathfrak{a}_{2},\mathfrak{h}_{2})$ reversed) this is the same as case (2).